# Accounting for motion of supernova host galaxy in statistical inference from SNIa data

### v1

@@ Abstract
We investigate the impact of peculiar motion of Type Ia supernova host galaxies on cosmological parameter estimation. This motion causes their redshift to deviate from that of the comoving observer at their position and is a source of noise. To this end, we develop an estimator for parameter estimation in models with errors in independent variables. Using the Bayesian framework, errors in independent variables are treated as nuisance parameters by making the independent variables parameters of the model. Our method applied to the Pantheon sample of Type Ia supernova indicates a few percent shift in the central values of inferred cosmological parameters. For the $w$CDM model, we find that accounting for peculiar velocities makes the data marginally more consistent with the cosmological constant model. By using simulated data, we show that not accounting for peculiar velocities will significantly impact parameter estimation from higher precision future data sets.
1 Introduction
@@
The spatially flat $\Lambda$CDM model, based on the cosmological principle of homogeneity and isotropy, constituting cold dark matter and the cosmological constant that dominates the universe at present, has become the standard description of our observed universe. It has been remarkably successful in describing the observed anisotropies in the Cosmic Microwave Background (CMB)[1], large-scale clustering of matter [2, 3, 4], and the background expansion history inferred from the redshift and brightness measurements of Type Ia supernovae (SNIa) [5, 6, 7]. In addition, a growing number of space and earth missions have ushered in the era of precision cosmology [8], which allows us to constrain cosmological parameters at a sub-percent level of precision. For example, the measurement of the CMB temperature and polarization anisotropies by the Planck mission has allowed us to constrain the Hubble constant within an uncertainty of $0.5$ km/s/Mpc [1]. This increasing quality and quantity of data has enabled us to extract more detailed information about the present and past of the universe.
@@
However, the onset of the era of greater precision has also revealed systematic discrepancies in the standard model of cosmology, such as the observed tensions in $H_{0}$ and $S_{8}$ [9, 10, 11]. Taking these discrepancies as a hint of new physics beyond the standard $\Lambda$CDM model, several modifications to the theory have been proposed in the literature, on the contrary, untreated or systematic errors in the analysis as a source of observed discrepancies remain a less explored but plausible alternative. For example, there are certain anomalies in the CMB observations, such as lensing anomalies, whose presence depends on the method used for statistical analysis. For example, the standard Plik and Commander likelihood of Planck 2018 data prefer a higher value of the lensing parameter $A_{L}$ than the prediction of the $\Lambda\rm CDM$ model, $A_{L}=1$. However, in the likelihood NPIPE based on the same Planck data but different modeling of systematics, there is no such anomaly [12]. It was shown in [13] that alleviating these tensions with extensions to the $\Lambda$CDM model is not sufficient to claim a hint for physics beyond the $\Lambda$CDM model. This also highlights the fact that increasing the precision in measurement also requires more accurate statistical analysis techniques to obtain consistent results from different observational probes.
@@
In statistics literature, the study of model fitting when there are errors in both dependent and independent variables is called errors-in-variables models [14]. Many different methods have been proposed to deal with this issue if additional information can be obtained about the independent variable. The instrumental variable method [14] considers another variable that is related to the true regressor with some known relation. The method of repeated measurement [14] uses multiple observations of the same variable to minimize error. For the special case, when the model is a monotonic and differentiable function of the regressor variable, the likelihood can be written by inverting the model function [15]. However, there is no general method for fitting arbitrary errors-in-variable models.
@@
In the context of cosmology, this issue could arise on various occasions. For example, it becomes relevant in estimating the cosmological parameters from the magnitude–redshift data of SNIa. Usually, the redshift determination of these supernovae is considered to be free of errors. There are at least two sources of errors in the determination of the redshift. Spectroscopic redshifts can be determined with a relative accuracy better than $10^{-4}$. The second source of error arises from the line-of-sight peculiar velocities owing to gravitational interaction [^1] Cosmological peculiar velocities are a key prediction of the gravitational collapse model of structure formation. Their impact can be detected in the redshift-space two-point correlation function of galaxies (e.g. [16] and the references therein for details). These velocities can also be detected statistically using the kSZ (Kinetic Sunayaev-Zeldovich) effect (e.g. [17, 18]). The line-of-sight peculiar motion of host galaxies of supernovae contributes to the total observed redshift and cannot be disentangled from the contribution from cosmological expansion [^2] The contribution of peculiar motion also arises from our motion with respect to the cosmic rest frame. The magnitude and direction of this component can be determined from the measurement of the CMB dipole. It is corrected for in the available SNIa data as in the Pantheon sample.. From linear perturbation theory, the RMS of peculiar velocity varies from a few hundred km/sec to a few thousand km/sec for scales 1 Mpc to 100 Mpc. The modeling of the peculiar motion of individual objects requires taking into account nonlinear effects, which could further enhance this effect. The impact of peculiar velocities could be more than an order of magnitude larger than the instrumental uncertainty but still small compared to Hubble recessional velocities except in the nearby universe. The impact of peculiar motion on the cosmological parameters has been studied in the literature using N-Body simulations [19, 20, 21, 22]. The peculiar velocity correction in the Pantheon+ SNIa sample was obtained in [23]. These studies focus mainly on the determination of the Hubble constant.
@@
In this work, we analyze the impact of peculiar velocities on the determination of cosmological parameters using SNIa data by studying the problem of fitting models with errors in both dependent and independent variables. We analyze the Pantheon SNIa sample [24] using three estimators—one without peculiar motions, one based on a locally linear magnitude–redshift relationship, and finally, one that accounts for the theoretical model without any approximation. We then apply our estimators on simulated data with parameters (e.g. number of SNIa, errors on magnitude) typical of the current and future SNIa data sets. For our study, we consider two cosmological models: a spatially flat $\Lambda$CDM model and a $w$CDM model with constant $w$.
@@
This paper is organized as follows. In Section 2, we discuss in general our method for dealing with the problem of fitting arbitrary models with errors in both variables. In Section 3, we present the magnitude–redshift relation as an errors-in-variables model. In Section 4, we apply this method to estimate cosmological parameters from supernova data for the $\Lambda$CDM and the $w$CDM model, and discuss the modification in the results as compared to the standard model fitting method that does not consider the contribution of peculiar motion. In Section 5, we conclude with a discussion of the results and potential future applications of our method. In Appendix A, we derive an expression for likelihood in the linear approximation of the magnitude–redshift relation.
2 Methodology
@@
The problem of regression with an arbitrary errors-in-variables model is interesting and, to the best of our knowledge, not a commonly addressed problem in astronomy. Since astronomical data usually have errors in independent variables, this could have general applications in a variety of problems in astronomy. In this section, we describe a Bayesian approach to this problem. In Section 2.1, we describe our method in general while motivating our choice of the particular form of estimator that we propose. Then, in Section 3, we specialize it for the estimation of cosmological parameters from the magnitude–redshift data of the Pantheon sample of supernovae [24]. We consider the contribution of peculiar motion of supernova host galaxies to the redshift uncertainty, which leads to its formulation as an errors-in-variables problem. We also compare it to an approach that uses a local linear approximation of the magnitude–redshift relation and uses known methods of handling errors-in-variables for linear models.
2.1 Estimator for Errors-in-Variables
@@
Quite often, the output of experiments is a pair $(x_{i},y_{i})$ [^3] It is worth noting that in general, the data could have several independent variables $(x_{i},y_{i},z_{i},\cdots)$. It is straightforward to generalize this analysis to such multivariate cases. of $N$ data points with uncertainties, assuming the randomness to be Gaussian, $\sigma_{x_{i}}$ and $\sigma_{y_{i}}$, associated with each observed value $x_{i}$ and $y_{i}$ respectively. The quantity $y_{i}$ often depends on the independently measured quantity $x_{i}$ through a model that could be given through a mathematical relation
$y_{i}=f(\theta,x_{i}),$ (2.1)
@@
or, in general, could be given by an algorithm such as a numerical simulation or a complex mathematical operation. Here, $f(\theta,x_{i})$ is a function of the $m$ model parameters, $\theta\equiv(\theta_{1},\theta_{2},...\theta_{m})$ and the variable $x_{i}$. In our notation, if any expression depends explicitly on a single variable, then we notate it with a subscript; else, we use unscripted symbols to refer to them collectively. We shall soon focus on the case of apparent magnitude versus redshift data that comprises the measured pair $(z_{i},m_{i})$, which might be helpful to the reader while reading the following discussion to make it less abstract.
@@
If the function $f(\theta,x_{j})$ is a linear function of $x_{j}$, then estimating the parameters $\theta$ from data that have errors on both axes is quite straightforward (see, for example, [25] and the derivation in Appendix A). This method is based on the conversion of the $x$ error into the $y$ error through the slope of the linear function and adding it in quadrature to the $y$ error while constructing the likelihood function. However, when the model is nonlinear in $x_{j}$, this trick works only by approximating the fitting function as locally a linear function, as shown for the special case of the magnitude–redshift relation in the Appendix A. Although, in general, a nonlinear model could refer to nonlinear dependence on $\theta$ or $x_{i}$, in this paper, by nonlinear model, we always refer to the latter case.
2.1.1 Estimator for nonlinear models
@@
To keep things general, let $P_{\rm Y}(\epsilon_{Y})$ be the probability distribution function for random statistical noise in the measurement of $y$, so that the observed values are offset from their true value by a random number, $y_{i}=f(\theta,x_{i})+\epsilon_{Yi}$, where $\epsilon_{Yi}$ is the random error in the measurement $y_{i}$. Therefore, the likelihood function is given by
$\mathcal{L}(\mathcal{D}\mid\theta)=\prod_{i}P_{Y}(y_{i}-f(\theta,x_{i})),$ (2.2)
@@
where $\mathcal{D}$ collectively refers to the pair $(x,y)$. In Bayesian statistics, using Bayes’ theorem, this likelihood function can be used to write the posterior probability distribution for parameters $\theta$ as
$P(\theta\mid\mathcal{D};I)\propto\mathcal{L}(\mathcal{D}\mid\theta;I)P(\theta\mid I)$ (2.3)
where $P(\theta\mid I)$ is the prior probability distribution function for $\theta$, and the addition of $I$ in various terms symbolizes all the prior information that we may have (for details, see, e.g., [26]). Until now, the variables $x$ have been considered given and fixed. To incorporate error in variables $x$ in the posterior distribution, we now turn them into undetermined parameters $x^{*}$, equivalent to parameters $\theta$, in terms of which the likelihood function can be written as
$\mathcal{L}(\mathcal{D}\mid\theta,x^{*};I)=\prod_{i}P_{Y}(y_{i}-f(\theta,x^{*}_{i}))$ (2.4)
Note that conversion of $x$ to parameter means that now the data $\mathcal{D}$ refers only to $y$. The information about errors in $x$ and their measured values can be brought into our analysis as a prior on $x^{*}$ through [^4] We can derive the same expressions by noting that the joint probability of errors $\epsilon_{X}$ and $\epsilon_{Y}$ is given by $P_{X}(\epsilon_{X})P_{Y}(\epsilon_{Y})$, thus the errors on $x$ can be considered to be a part of likelihood of a data set comprising the pair $(x,y)$
$P(x^{*}_{i}\mid I)=P_{\rm X}(x^{*}_{i}-x_{i}),$
where $P_{\rm X}$ is the probability distribution for errors in variable $x$ (assumed for simplicity to be the same for all data points, but this can be easily generalized). The modified posterior distribution is given by
$P(\theta,x^{*}\mid\mathcal{D};I)\propto\prod_{i}P_{Y}(y_{i}-f(\theta,x^{*}_{i}))P(x^{*}_{i}\mid I)P(\theta\mid I)$ (2.5)
If the random noise is Gaussian, then the log of the likelihood function is given by
$\ln\mathcal{L}(\mathcal{D}\mid\theta,x^{*})=-\frac{1}{2}\sum_{i=1}^{N}\left[\frac{y_{i}-f(\theta,x^{*}_{i})}{\sigma_{y_{i}}}\right]^{2}\equiv-\frac{\chi^{2}_{Y}(\theta,x^{*})}{2},$ (2.6)
and the log of the prior probability distribution function for $x$ is given by
$\ln{P(x^{*}_{i}\mid I)}=-\frac{1}{2}\sum_{i=1}^{N}\left[\frac{x^{*}_{i}-x_{i}}{\sigma_{x_{i}}}\right]^{2}\equiv-\frac{\chi^{2}_{X}(x^{*})}{2}$ (2.7)
In terms of these variables, the posterior distribution can be written as
$P(\theta,x^{*}\mid\mathcal{D})\propto\exp\left[-\frac{\chi^{2}_{Y}(\bm{\theta},x^{*})+\chi^{2}_{X}(x^{*})}{2}\right]P(\bm{\theta}\mid I)$ (2.8)
We can marginalize $x^{*}$ to obtain the probability distribution for $\theta$
$P(\theta\mid\mathcal{D})\propto\int d^{N}x^{*}\exp\left[-\frac{\chi^{2}_{Y}(\bm{\theta},x^{*})+\chi^{2}_{X}(x^{*})}{2}\right]P(\bm{\theta}\mid I)$ (2.9)
In practice, marginalization is more conveniently carried out through MCMC, as we discuss later.
@@
Since the total number of parameters now exceeds the number of data points, there is, in principle, the possibility of overfitting the data. In order to obviate this possibility, we add a regularization term as a prior probability that our physical model should be such that the mean value of $\chi_{Y}^{2}$ is close to degrees of freedom of the model [^5] This expectation is technically true only for linear models, but holds quite well in general for small errors as models can then be roughly treated as locally linear.. For the case of Gaussian noise in the dependent variable, this is just the probability distribution function for $\chi_{Y}^{2}$ given by
$P_{\chi^{2}}(\chi^{2}_{Y})=\frac{\chi_{Y}^{k-2}\exp(-\chi^{2}_{Y}/2)}{2^{k/2}\Gamma(k/2)},$ (2.10)
@@
where $k=N-m$ is the desired degrees of freedom of the fit (not accounting for noise in the $x$ values). For example, for the case of fitting luminosity distance to distant supernovae, this would be the total number of supernovae minus the number of cosmological parameters being fitted. Therefore, our final form for the posterior distribution function for $\theta$ can be obtained by multiplying it with the probability distribution function for $\chi_{Y}^{2}$
$P(\theta\mid\mathcal{D})\propto\int d^{N}x^{*}\exp\left[-\frac{\chi^{2}_{Y}(\bm{\theta},x^{*})+\chi^{2}_{X}(x^{*})}{2}\right]P(\bm{\theta}\mid I)P_{\chi^{2}}(\chi_{Y}^{2})\,.$ (2.11)

### v2

@@ Abstract
We introduce a Bayesian method to estimate peculiar velocities of Type Ia supernova (SNIa) host galaxies by employing the magnitude-redshift relationship of SNIa. Random peculiar motions act as noise in the estimation of redshift, and constitute independent variables in the SNIa data. We develop a method to take into account errors in independent variables for general nonlinear models. Using the MCMC sampling technique, we implement numerical codes to estimate peculiar velocities along with other cosmological parameters. In this paper, we study the impact of these velocities on cosmological parameters by treating them as nuisance parameters. We apply our proposed method on the Pantheon sample of SNIa and show a few percent shift on the central values of inferred cosmological parameters. Although the current data are not particularly sensitive to this error, using simulated data, we also gauge the efficacy of our method on the future SNIa observations and show that future data require the inclusion of this error in cosmological parameter estimation. Our method complements several existing methods that seek to estimate peculiar velocities using galaxy data and N-body simulations and extends to higher redshifts.
Cosmology – Bayesian statistics, Peculiar motion, Type Ia supernovae.
I Introduction
@@
The spatially flat $\Lambda$CDM model, based on the cosmological principle of homogeneity and isotropy, constituting cold dark matter and the cosmological constant that dominates the universe at present, has become the standard description of our observed universe. It has been remarkably successful in describing the observed anisotropies in the Cosmic Microwave Background (CMB)Aghanim et al. (2020), large-scale clustering of matter Kwan et al. (2017); Cole et al. (2005); Adame et al. (2024), and the background expansion history inferred from the redshift and brightness measurements of Type Ia supernovae (SNIa) Conley et al. (2010); Brout et al. (2022); Scolnic et al. (2022). In addition, a growing number of space and earth missions have ushered in the era of precision cosmology Turner (2022), which allows us to constrain cosmological parameters at a sub-percent level of precision. For example, the measurement of the CMB temperature and polarization anisotropies by the Planck mission has allowed us to constrain the Hubble constant within an uncertainty of $0.5$ km/s/Mpc Aghanim et al. (2020). This increasing quality and quantity of data has enabled us to extract more detailed information about the present and past of the universe.
@@
However, the onset of the era of greater precision has also revealed systematic discrepancies in the standard model of cosmology, such as the observed tensions in $H_{0}$ and $S_{8}$ Di Valentino et al. (2021a, b); Abdalla et al. (2022). Taking these discrepancies as a hint of new physics beyond the standard $\Lambda$CDM model, several modifications to the theory have been proposed in the literature. However, untreated or systematic errors in the analysis as a source of observed discrepancies remain a less explored but plausible alternative. For instance, there are certain anomalies in the CMB observations, such as lensing anomalies, whose presence depends on the method used for statistical analysis. Specifically, the standard Plik and Commander likelihood of Planck 2018 data prefer a higher value of the lensing parameter $A_{L}$ than the prediction of the $\Lambda\rm CDM$ model, $A_{L}=1$. However, in the likelihood NPIPE based on the same Planck data but different modeling of systematics, there is no such anomaly Rosenberg et al. (2022). It was argued in Cortês and Liddle (2024) that alleviating these tensions with extensions to the $\Lambda$CDM model is not sufficient to claim a hint for physics beyond the $\Lambda$CDM model. This also highlights the fact that increasing the precision in measurement also requires more accurate statistical analysis techniques to obtain consistent results from different observational probes.
@@
In statistics literature, the study of model fitting when there are errors in both dependent and independent variables is called errors-in-variables models Dougherty (2007). Many different methods have been proposed to deal with this issue if additional information can be obtained about the independent variable. The instrumental variable method Dougherty (2007) considers another variable that is related to the true regressor with some known relation. The method of repeated measurement Dougherty (2007) uses multiple observations of the same variable to minimize error. For the special case, when the model is a monotonic and differentiable function of the regressor variable, the likelihood can be written by inverting the model function Gregory (2005). However, there is no general method for fitting arbitrary errors-in-variable models.
@@
In the context of cosmology, this issue could arise on various occasions. For example, it becomes relevant in estimating the cosmological parameters from the magnitude–redshift data of SNIa. Usually, the redshift determination of these supernovae is considered to be free of errors. There are at least two sources of errors in the determination of the redshift. Spectroscopic redshifts can be determined with a relative accuracy better than $10^{-4}$. The second source of error arises from the line-of-sight peculiar velocities owing to gravitational interaction [^1] Cosmological peculiar velocities are a key prediction of the gravitational collapse model of structure formation. Their impact can be detected in the redshift-space two-point correlation function of galaxies (e.g. Dodelson and Schmidt (2020) and the references therein for details). These velocities can also be detected statistically using the kSZ (Kinetic Sunayaev-Zeldovich) effect (e.g. Aghanim et al. (2018); Tanimura et al. (2021)). The line-of-sight peculiar motion of host galaxies of supernovae contributes to the total observed redshift and cannot be disentangled from the contribution from cosmological expansion [^2] The contribution of peculiar motion also arises from our motion with respect to the cosmic rest frame. The magnitude and direction of this component can be determined by measuring the CMB dipole. It is corrected for in the available SNIa data as in the Pantheon sample.. From linear perturbation theory, the RMS of peculiar velocity varies from a few hundred km/sec to a few thousand km/sec for scales 1 Mpc to 100 Mpc. The modeling of the peculiar motion of individual objects requires taking into account nonlinear effects, which could further enhance this effect. The impact of peculiar velocities could be more than an order of magnitude larger than the instrumental uncertainty but still small compared to Hubble recessional velocities except in the nearby universe.
The impact of peculiar motions on the estimation of cosmological parameters using SNIa data has been studied in the literature using N-Body simulations (e.g. Gavas et al. (2024); Wu and Huterer (2017); Wojtak et al. (2014); Peterson et al. (2024)) and galaxy data (e.g. Davis et al. (2011); Peterson et al. (2022); Giani et al. (2024)). In particular, Peterson et al. (2022) modeled peculiar velocities within the framework of the $\Lambda$CDM model as a sum of a large-scale correlated component and a small-scale random component and estimated them independently of the SNIa data using complementary astrophysical probes and finally estimated their impact on the Hubble parameter. Using a frequentist technique, Davis et al. (2011) modeled the peculiar velocity variance as an additional term in the covariance matrix for the magnitude errors of SNIa to compute the corrections these velocities induce in the estimation of cosmological parameters such as the Hubble constant $H_{0}$ and the dark energy equation of state parameter $w$ from the SNIa data. This method is somewhat similar to one of our estimators based on an approximate locally linear magnitude–redshift relationship discussed in Section IV. However, our estimator uses the slope of the magnitude-redshift relationship to propagate redshift errors into the likelihood function, while Davis et al. (2011) used a constant term.
Our main approach to accounting for peculiar velocities in the SNIa data is more direct and makes minimal assumptions about the origin of this motion. Furthermore, our method does not require additional input from independent astrophysical probes of peculiar motion. It is based on treating each velocity as an independent parameter drawn from a flat prior centered at the position of the measured redshift, which means that no specific underlying cosmological model is assumed for peculiar motion. Each velocity is estimated using a Bayesian method based on likelihood function and then marginalized to assess their impact on cosmological parameters. To this end, we develop a Bayesian method that is applicable to the problem of fitting general non-linear models with errors in both dependent and independent variables.
@@
We analyze the Pantheon SNIa sample Scolnic et al. (2018) using three estimators: 1) one without including peculiar motions, 2) one based on a locally linear magnitude–redshift relationship, with an additional assumption of Gaussian errors in redshift, and finally 3) one that accounts for peculiar motions without making any approximation. The second estimator was considered for comparison with the more general third estimator, as it is computationally much less expensive. To evaluate the efficacy of our estimators, we also applied our estimators to synthetic data with parameters (e.g., number of SNIa, errors on magnitude) typical of the current and future SNIa data sets. This is useful to establish the efficacy of our estimators for the current quality of data and also to demonstrate that higher-quality future data sets would lead to incorrect estimates of cosmological parameters without proper accounting of peculiar motions. For our study, in all cases we consider two cosmological models: a spatially flat $\Lambda$CDM model and a $w$CDM model with constant $w$.
@@
This paper is organized as follows. In Section II, we discuss in general our method for dealing with the problem of fitting arbitrary models with errors in both variables. In Section III, we present the magnitude–redshift relation as an error-in-variables model. In Section IV, we apply this method to estimate cosmological parameters from supernova data for the $\Lambda$CDM and the $w$CDM model, and discuss the modification in the results as compared to the standard model fitting method that does not consider the contribution of peculiar motion. In Section IV, we conclude with a discussion of the results and potential future applications of our method. In Appendix A, we derive an expression for likelihood in the linear approximation of the magnitude–redshift relation.
II Methodology
@@
The problem of regression with an arbitrary error-in-variables model is an interesting problem and, to the best of our knowledge, not a commonly addressed problem in astronomy. Since astronomical data usually have errors in independent variables, this could have general applications in a variety of problems in astronomy. In this section, we describe a Bayesian solution to this problem. In Section II.1, we begin by describing the general background of the problem and develop our estimator by motivating it with the familiar case of independent variables without errors. Then, in Section III, we specialize our estimator for estimating cosmological parameters from the magnitude–redshift data of the Pantheon sample of supernovae Scolnic et al. (2018). We consider the contribution of peculiar motion of supernova host galaxies to the redshift uncertainty, which leads to its formulation as an errors-in-variables problem. Later in this paper, we have also compared it to an approximate method that uses a local linear approximation of the magnitude–redshift relation and assumes a Gaussian distribution of errors due to peculiar motions. The development of this method can be found in the Appendix A. This comparison is important because the approximate method is computationally inexpensive and is therefore useful for very large data sets. Our results show that, despite its assumptions, it gives comparable results to the exact method. However, the exact estimator is more flexible in handling a greater variety of problems.
II.1 Estimator for Errors-in-Variables
@@
Frequently, experiments produce $N$ pairs $(x_{i},y_{i})$ of noisy data points with uncertainties (assuming the randomness to be Gaussian) $\sigma_{x_{i}}$ and $\sigma_{y_{i}}$ associated with each observed value $x_{i}$ and $y_{i}$, respectively. The quantity $y_{i}$ often depends on the independently measured quantity $x_{i}$ through a physical model that is often given in terms of a mathematical relation,
$y_{i}=f(\theta,x_{i}),$ (1)
@@
or, in general, could be given by an algorithm such as a numerical code or a multistep mathematical operation that outputs $y$ for every $x$. However, our treatment is based on expressing the association of each $x$ with $y$ abstractly through the above relationship. The function $f(\theta,x_{i})$ is generally a function of $m$ unknown model parameters $\theta\equiv(\theta_{1},\theta_{2},...\theta_{m})$ along with the variable $x_{i}$. In the notation used in this paper, if any expression depends explicitly on a single variable, then we notate it with a subscript; else, we use unscripted symbols to refer to them collectively. We shall soon focus on the case of apparent magnitude-redshift data that comprise the measured pair $(z_{i},m_{i})$. It might be helpful to the reader to keep this concrete model in mind while reading this section. The parameters $\theta$ would then be the usual cosmological parameters such as the Hubble constant, various matter energy densities, and the equation of state parameter.
@@
The most straightforward case is where the independent variable $x$ can be measured with negligible error. This is often how such data are fitted to models, often ignoring the $x$ errors even if they are not small. This is due to the fact that the canonical method relies on working with the likelihood function derived from the errors in the dependent variable $y$. Therefore, we must find a way to incorporate $x$ errors as an additional source of error in the likelihood function. If the function $f(\theta,x_{j})$ is a linear function [^3] Although in general a nonlinear model could refer to a model $f(\theta,x_{j})$ with nonlinear dependence on $\theta$ or $x_{i}$, in this paper by a nonlinear model we always refer to the latter case, of $x_{j}$, then estimating the parameters $\theta$ from data that have errors in both variables is well established and quite straightforward (see, for example, Press et al. (1992) and the derivation in Appendix A). This method is based on the conversion of the $x$ error into the $y$ error through the slope of the linear function and adding it in quadrature to the $y$ error while constructing the likelihood function. However, when the model is nonlinear in $x_{j}$, this trick works only by approximating the fitting function as locally a linear function, as shown for the special case of the magnitude–redshift relation in the Appendix A. Our aim in this paper is to develop a method that is not approximate or ad hoc and treats errors in $x$ and $y$ on equal footing. This turns out to be possible in Bayesian statistics, as we show next.
II.1.1 Estimator for nonlinear models
@@
We introduce our approach through the familiar case of data with errors only in the dependent variable $y$ and none in the independent variable $x$. Let the true values of $y$ be denoted by $y^{*}$. The observed values of $y$ differ from their true value by a random number, $y_{i}=y^{*}_{i}+\epsilon_{Yi}$, where $\epsilon_{Yi}$ is the random error in measurement $y_{i}$. Since $x$ are measured without error, $y^{*}_{i}=f(\theta,x_{i})$. If $P_{\rm Y}(\epsilon_{Y})$ is the probability distribution function for random noise in the measurement of $y$, the likelihood function is given by
$\mathcal{L}_{Y}(\mathcal{D}\mid\theta,y^{*})=\prod_{i}P_{Y}(y_{i}-y^{*}_{i})\delta(y^{*}_{i}-f(\theta,x_{i})),$ (2)
where $\mathcal{D}$ collectively refers to the pair $(x,y)$. In this form, the likelihood function can only be evaluated if we know the values of $\theta$ and $y^{*}$. The Dirac delta function ensures that the likelihood is non-zero only when $y^{*}_{i}=f(\theta,x_{i})$. Since $x_{i}$ is known but not $y^{*}_{i}$, this equation can be satisfied for any value of $y^{*}_{i}$ at $\theta$ differing from their true value. Furthermore, the subscript $Y$ in the likelihood function has been added to make it explicit that it is derived from the random noise in $y$. In a sense, we can think of the true values $y^{*}$ as additional parameters of the model, which is how they are treated in the following analysis.
@@
Using Bayes’ theorem, this likelihood function can be used to write the posterior probability distribution for the parameters $\theta$, and $y^{*}$ as
$P(\theta,y^{*}\mid\mathcal{D};I)\propto\mathcal{L}_{Y}(\mathcal{D}\mid\theta,y^{*})P(\theta\mid I)$ (3)
where $P(\theta\mid I)$ is the prior probability distribution function for $\theta$, and the addition of $I$ in various terms symbolizes all the prior information that we may have (for details, see, e.g., Sivia (1996)). We also assumed that there is no prior information on the true values $y^{*}$. Combining Eq. (2) with this equation we obtain
$P(\theta,y^{*}\mid\mathcal{D};I)\propto\prod_{i}P_{Y}(y_{i}-y^{*}_{i})\delta(y^{*}_{i}-f(\theta,x_{i}))P(\theta\mid I)\,.$ (4)
This posterior probability distribution has information about the true values $y^{*}$, which is of little use. Usually, the parameters $\theta$ are of interest; therefore, we can marginalize over $y^{*}$ by integrating over them to obtain the most commonly used expression for the posterior probability for a model with errors only in dependent variable,
$P(\theta\mid\mathcal{D};I)\propto\prod_{i}P_{Y}(y_{i}-f(\theta,x_{i}))P(\theta\mid I)\,.$ (5)
To generalize this to the case with errors in both variables, in analogy with $y$ we write $x_{i}=x^{*}_{i}+\epsilon_{Xi}$, where $x^{*}$ is the unknown true value of $x$ in terms of which the likelihood function can be written as
$\mathcal{L}_{XY}(\mathcal{D}\mid\theta,x^{*},y^{*})=\prod_{i}P_{Y}(y_{i}-y^{*}_{i})P_{X}(x_{i}-x^{*}_{i})\delta(y^{*}_{i}-f(\theta,x^{*}_{i})),$ (6)
where we have used the fact that $x$ and $y$ are statistically independent, therefore, their joint probability can be obtained simply by multiplying their individual probability distribution functions. However, note that since $x$ are now noisy, the argument of delta function now contains only true values $(x^{*},y^{*})$. The modified posterior distribution is now given by
$P(\theta,x^{*},y^{*}\mid\mathcal{D};I)\propto\prod_{i}P_{Y}(y_{i}-y^{*}_{i})P_{X}(x_{i}-x^{*}_{i})\delta(h(x^{*},y^{*};\theta))P(\theta\mid I)$ (7)
where we have defined $h(x^{*},y^{*};\theta)\equiv y^{*}_{i}-f(\theta,x^{*}_{i})=0$ as a symmetric way to write the model relationship between $x$ and $y$. The obvious advantage is that it makes the difference between dependent and independent variables disappear. It is worth noting that in general, the data could have several independent variables $(x_{i},y_{i},z_{i},\cdots)$ related through $h(x_{i},y_{i},z_{i},\cdots)=0$. It is straightforward to generalize this analysis to such multivariate cases.
The posterior distribution in Eq. (7) is a function not only of the model parameters but also of the unknown true values of $x$ and $y$ for each data point. If we are interested only in the model parameters $\theta$, then we can integrate over additional parameters. This yields
$P(\theta\mid\mathcal{D};I)\propto\int d^{N}x^{*}\prod_{i}P_{Y}(y_{i}-f(\theta,x^{*}_{i}))P_{X}(x_{i}-x^{*}_{i})P(\theta\mid I)$ (8)
If we assume Gaussian errors in both variables with standard deviations $\sigma_{y}$ and $\sigma_{x}$, respectively, then after defining
$\chi^{2}_{Y}(\theta,x^{*})=\sum_{i=1}^{N}\left[\frac{y_{i}-f(\theta,x^{*}_{i})}{\sigma_{y_{i}}}\right]^{2},\quad\chi^{2}_{X}(x^{*})=\sum_{i=1}^{N}\left[\frac{x^{*}_{i}-x_{i}}{\sigma_{x_{i}}}\right]^{2}\,,$ (9)
the marginalized posterior distribution can be written as
$P(\theta\mid\mathcal{D};I)\propto\int d^{N}x^{*}\exp\left[-\frac{\chi^{2}_{Y}(\theta,x^{*})+\chi^{2}_{X}(x^{*})}{2}\right]P(\theta\mid I)\,.$ (10)
This form is used in Appendix A to derive an approximate form of the estimator where integration over $x^{*}$ can be done analytically. However, if we do not assume a specific form for $P_{X}$ then assuming Gaussian errors only in $y$ we get
$P(\theta\mid\mathcal{D};I)\propto\int d^{N}x^{*}\exp\left[-\frac{\chi^{2}_{Y}(\theta,x^{*})}{2}\right]P_{X}(x_{i}-x^{*}_{i})P(\theta\mid I)\,.$ (11)
@@
In practice, the integration in Eq. (11) is more conveniently carried out with Markov Chain Monte Carlo (MCMC), as we discuss later. Since the total number of model parameters now exceeds the number of data points, there is, in principle, the possibility of overfitting the data. In order to obviate this possibility, we add a regularization term to ensure that the physical model satisfy the condition that the mean value of $\chi_{Y}^{2}$ is close to the degrees of freedom of the model [^4] This expectation is technically true only for linear models, but holds well in general for small errors as models can then be roughly thought of as locally linear.. The marginalized $x^{*}$ can be considered as subsidiary parameters not directly related to the physical model. Therefore, the degrees of freedom is given by the number of data points, $N$, minus the number of physical parameters, $m$. For the case of Gaussian noise in the dependent variable, the required regularization term is just the probability distribution function for $\chi_{Y}^{2}$ given by
$P_{\chi^{2}}(\chi^{2}_{Y})=\frac{\chi_{Y}^{k-2}\exp(-\chi^{2}_{Y}/2)}{2^{k/2}\Gamma(k/2)},$ (12)
@@
where $k=N-m$ is the desired degrees of freedom of the fit. For the case of fitting magnitude-redshift relationship for distant supernovae, this would be the total number of supernovae minus the number of cosmological parameters being fitted. Therefore, our final form for the posterior distribution function for $\theta$ can be obtained by multiplying it with the probability distribution function for $\chi_{Y}^{2}$
$P(\theta\mid\mathcal{D};I)\propto\int d^{N}x^{*}\exp\left[-\frac{\chi^{2}_{Y}(\theta,x^{*})}{2}\right]P_{X}(x_{i}-x^{*}_{i})P({\theta}\mid I)P_{\chi^{2}}(\chi_{Y}^{2})\,.$ (13)